### v1

@@ VI-B LLM-aided Reinforcement Learning for Network Optimization
Reinforcement learning is one of the most important techniques for network optimization. It explores various sequential action combinations, e.g., network resource allocation strategies and signal transmission power level, to maximize the long-term reward, such as higher data rate or lower transmission delay[168]. Many network optimization problems can be transformed into a unified MDP, and then using reinforcement learning to improve network metrics dynamically. For instance, resource allocation is a very common problem in many telecom scenarios, in which allocation decisions, desired network performance metrics, and network dynamics are usually defined as actions, rewards, and states, respectively[6]. However, it is worth noting that these definitions are usually intuitive and require expert knowledge of reinforcement learning techniques and telecom. Especially, most reward functions are manually designed using trial-and-error approaches, and the algorithm performance is affected by the hyperparameter selection, e.g., learning rate, batch size, and number of hidden layers. Fortunately, LLM techniques provide new opportunities to overcome these bottlenecks. This section will introduce two LLM-aided reinforcement learning techniques: automatic reward function design and verbal reinforcement learning.
@@ VI-B1 Using LLM for reward function design
A recent survey in [27] shows that 92% reinforcement learning researchers use manual trial-and-error reward function design and 89% indicate that the designed rewards lead to unintended behaviour during algorithm training[169]. Such issues become more difficult in complicated telecom scenarios since various network elements are involved, e.g., users with diverse requirements, limited available resources, and dynamic network environments. To this end, LLM shows the capability of developing a universal approach for reward design, which will significantly lower the difficulty of using reinforcement learning. For instance, [34] proposed a self-refined LLM model for automated reward function design in robotics, achieving a comparable performance as manually designed functions. [35] applies LLM as a proxy reward function, where the user provides a textual prompt with a few examples or a description of the desired behaviour. In the following, we will introduce how the automatic reward function is designed.
@@
  • A mapping function $\mathcal{M}$ that maps the textual output of LLM to a binary value, e.g., ”good” or ”bad”. This binary value feedback indicates the quality criteria of the generation, and then the LLM model easily understands the overall feedback.
Given these definitions, the LLM-aided MDP definition becomes $<S,A,R,T,\mathcal{L},\mathcal{M}>$, where $\mathcal{L}$ is a set of prompts $l_{1}$ to $l_{4}$, and $\mathcal{M}$ is the mapping function. The reward function $\mathcal{F}$ in $R=\mathcal{F}(s,a)$ is defined by $\mathcal{F}:=\mathcal{G}(\mathcal{L},\mathcal{M})$, where $\mathcal{G}$ is the inference of LLM models. $\mathcal{F}:=\mathcal{G}(\mathcal{L},\mathcal{M})$ shows that the design of the reward function $\mathcal{F}$ depends on prompt input $\mathcal{L}$ and the mapping function $\mathcal{M}$. Based on the LLM-aided MDP framework, using LLM for reward function design can be summarized as the following steps:
  • Step 1: Description input. Using language to describe the task, objective, states, and actions. If necessary, providing possible trajectory examples to LLM models. Here an alternative approach is to feed the environment code to the LLM agent, and then using natural language to describe the task, which has been used in [36].
@@
  • Step 4: Evaluation and feedback. Evaluating the reinforcement learning training output and providing feedback to LLM models.
@@ VI-B2 Verbal reinforcement learning via LLM
  • Actor: The actor is built upon an LLM model, which is specifically prompted to generate actions, e.g., network control and optimization strategies. Based on short-term and long-term memories, the actor can apply various methods to produce actions, such as CoT [33] and ReAct [171]. These advanced prompt techniques can improve the actor’s capability of reasoning and planning, which can better adapt to the complicated decision-making of network optimization problems.
@@
Compared with conventional reinforcement learning, the LLM-aided verbal learning technique has multiple advantages for telecom optimization: 1) Lowering the difficulty of implementing network optimization. Verbal reinforcement learning avoids the difficulty of tuning hyperparameters like learning rate, batch size, and training frequency. This will significantly lower the difficulty of applying artificial intelligence to network optimization. 2) Allowing for language instructions to guide network optimization policies. Specifically, the LLM-aided system allows for language instructions to guide the agent exploration, which is much more efficient than existing strategies such as $\epsilon$-greedy policy. Experienced network operators can provide language instructions to the LLM models directly, and no ML knowledge is required. 3) Reasoning and interpretable explanations for algorithm performance. One crucial advantage of LLM-aided systems is that they provide interpretable explanations of the algorithm and telecom system performance, and these experiences can further help understand network management policies.
Despite the advantages, LLM-aided reinforcement learning is still at a very early stage, and there are very few studies that apply this technique to the telecom field. In addition, specific telecom domain knowledge may be required to let the LLM better understand user demand. Therefore, professional wireless knowledge datasets such as TeleQnA in [110] may be required to fine-tune the LLM model.
@@ VI-D LLM-aided Convex Optimization for Telecom
Figure 15: LLM-aided meta-heuristic algorithm generation.
@@
  • Telecom knowledge and formulation templates: Telecom optimization requires professional network knowledge. LLM has learned fundamental knowledge in the pre-training period such as calculating information capacity using Shannon’s formula. However, using state-of-the-art telecom knowledge and formulation templates to fine-tune the LLM model can better improve the modelling accuracy. For instance, a dataset named TeleQnA is defined in [110], and it includes nearly 10000 communication field questions from both standards and research articles.
@@ VI-E LLM-based Heuristic Algorithm Design
Heuristic algorithms are very useful techniques for network management and optimization. Specifically, they apply diverse heuristic rules to select near-optimal solutions with low design and computational complexity[177]. Heuristic algorithms are particularly useful for solving optimization problems with integer control variables, which are very frequently formulated in telecom. For instance, the phase-shift optimization of RISs is considered as a very difficult problem with integer control variables and large solution space, and genetic algorithms and particle swarm optimization are used in [178] and [179] to solve this problem. In addition, heuristic algorithms are intuitively compatible with LLM models, since heuristic rules can be easily described by natural language and instructions. For example, swarm-based methods are very widely used heuristic algorithms, e.g., genetic algorithm, particle swarm optimization, and grey wolf optimizer, providing near-optimal solutions by iteratively searching for better solutions. However, the number of these algorithms has grown significantly in the past decade, and selecting the proper algorithm to solve specific network optimization problems has become more difficult. Given the reasoning and understanding capabilities, LLM offers promising changes for selecting and designing novel meta-heuristic algorithms.
@@ VI-F Discussions and Analyses
In addition, balancing LLM’s self-improvement and exploration-exploitation efficiency. It is straightforward that most previous techniques rely on the self-improvement capability of LLM models. For example, in verbal reinforcement learning, based on environmental feedback, the LLM agent can adjust the action selections to obtain a higher reward. On the other hand, exploration-exploitation difficulty is a common obstacle for many optimization techniques. This problem becomes severe when the action space becomes larger, which is very common in telecom networks. [36] proposed a solution by generating multiple solutions in each iteration and then applying evolutionary strategies to select better strategies. [159] showed a similar idea by using previous good trajectories as parents to generate new solutions. Therefore, how to use the LLM’s self-improvement capability and meanwhile balance the exploration-exploitation is very important.
Finally, although LLM already has fundamental knowledge of wireless networks in the pre-training period, fine-tuning LLM models with specific domain knowledge can still benefit telecom applications. In particular, a telecom-specific LLM model can better handle the domain-specific description, prompt, tasks, and solutions. For example, in reward function design and convex optimization problems in Sections VI-B1 and VI-D, we can directly use network terminologies to define the optimization problem, and the LLM can take full advantage of the domain knowledge to understand the questions.
@@ VII-A Motivations
Firstly, LLM models provide a universal and generalizable model for telecom network prediction. Given historical data, conventional prediction approaches must train a new model to adapt to incoming target tasks. These methods usually require extensive feature engineering and manual tuning, which can be time-consuming and may not generalize well across different scenarios. By contrast, the versatility of LLMs makes them suitable for processing diverse forms of time-series data, and such adaptability is crucial given the vast volumes of data generated in telecom. Moreover, LLM’s capability to continuously learn and adapt to new data patterns helps mitigate the concept drift problem, ensuring that the models remain relevant and effective over time. As a result, the integration of LLM techniques in time-series prediction offers a promising avenue for developing more robust and generalizable models that can better handle the complexities of data in telecom.
Meanwhile, LLM models have excellent ICL capabilities, which means that they can perform new tasks by leveraging contextual information in demonstrations. In particular, it means that the LLM can directly learn from the provided examples, and map the input-output relationships without extra model training. Such a prediction method is much more efficient than conventional prediction methods. Meanwhile, it is also more accessible since no professional knowledge of model training/fine-tuning is required. In addition, multi-modal LLM-enabled prediction can also be combined with sensing in telecom networks. In particular, multi-modal LLM models can process and integrate information from various data types, such as text, images, audio, and time-series data. In addition, sensing is an important part of envisioned 6G networks, aiming to integrate environmental information into communication networks, e.g., the image captured by street cameras or satellites, 3D LiDAR maps and WiFi sensing. In the context of telecom prediction, multi-modal LLMs can combine sensing data with numerical time-series data to generate more accurate context-aware prediction, which can be particularly useful in future 6G networks.
@@ VII-B Pre-training Foundation Models for Zero-shot Prediction
The pursuit of training a general-purpose foundation model for time-series data is driven by the desire to address the inherent challenges associated with diverse and dynamic data. Traditional time-series methods may struggle to adapt to the non-stationary properties of real-world data, where the statistical characteristics of the series can change over time due to evolving patterns and trends. For instance, the network traffic load level can be affected by many factors, including time, area, environment buildings, service types, etc. It usually requires dedicated model design and training for each target task, and then extracts the underlying patterns from history datasets[182]. By contrast, a general-purpose foundation model aims to overcome these challenges by leveraging the advancements in LLM technologies. The following will first formulate the problem of training a foundational LLM model for time-series prediction, and then discuss different tokenization mechanisms and model architectures.
@@ VII-B3 Model architecture
Decoder-only: Even though encoder-decoder models exhibit impressive effectiveness for handling sequences, decoder-only models become more popular in recent years. The essential difference between encoder-decoder and decoder-only models is that the bidirectional attention is used by encoder [41], which means each token is attending to all other tokens. In contrast, the decoder-only model employs casual attention, where each token cannot attend to tokens after it, but can only look at tokens before it. As shown by Fig.17, TimesFM [184] employs decoder-only architecture to train a time-series prediction model. Unlike traditional LLM techniques, which predict one element at a time, TimesFM is designed to predict extended future sequences in a single step, enhancing accuracy for long-term predictions. This flexibility also extends to inference; given a series, the model can predict its immediate future in fewer steps than a model with equal-length input and output segments would require. Such fast inference could be an appealing feature for telecom applications, because many prediction tasks require rapid response to network dynamics, such as channel state, short-term traffic changes, and indoor user locations. Conventional prediction methods usually take a long training time to adapt to such network environment changes. By contrast, TimesFM has the potential to capture short-term patterns instantly, which aligns with the fast-response requirements of telecom networks.
@@
With a backbone model and an effective tokenization mechanism, one can train a time-series prediction foundation model for telecom with a mixture of different datasets. Meanwhile, understanding the tokenization and model architecture differences is crucial for designing and pre-training a time series LLM for telecom applications. For instance, the tokenization mechanism introduced in the previous Section VII-B2 is very useful for telecom applications, since telecom networks are associated with a large number of network devices and end users, generating a huge number of datasets, such as historical CSI, traffic load level[185], and network performance metrics[186]. Therefore, reducing the number of input tokens can lower the pre-training difficulty of LLM models, especially considering that network edge devices usually have limited computational resources.
@@ VII-C Frozen Pre-trained LLM for Prediction
Rather than developing a specific LLM model for prediction, frozen pre-trained LLM refers to approaches that directly adapt a general-domain LLM to prediction tasks. This section delves into using a pre-trained LLM for prediction tasks without the necessity for further fine-tuning. There are two primary approaches: prompting-based and preprocessing-based methods. Specifically, the prompting-based methods include hard and soft prompts. Hard prompts employ rigid and predefined textual structures to present time-series information in a format that is intuitive for the language model. Conversely, soft prompts adopt a more nuanced strategy by integrating trainable embeddings within the input that subtly guide the language model’s predictions. Meanwhile, preprocessing-based methods aim to reformat the time series numerical values into a representation that aligns more seamlessly with LLM’s tokenization process, rather than introducing extra template tokens or trainable embeddings.
@@
Hard prompts ($P_{\text{hard}}$) involve pre-pending a fixed textual instruction or query to the input data sequence. Specifically, the model input for a time-series $x_{[1:T]}$ with a hard prompt is thus formalized as the concatenation $[P_{\text{hard}};x_{[1:T]}]$, which directs the model to generate a prediction in response to the prompt. When designing hard prompts for time-series prediction with language models, the general guideline is to transform numerical data into a format that mimics natural language constructs [187]. This involves two main components: input prompts and output prompts. Input prompts provide historical context and highlight the target time step for prediction, while output prompts focus on the desired prediction value, serving as the ground truth label for training or evaluation. Table X presents several telecom examples of designing hard prompts for specific tasks, such as network traffic load prediction, network user number prediction, and customer service prediction. The process mirrors the source/target structure common in machine translation tasks or can be likened to a question-answering setting, with the context as background information and the question seeking future insights. Then the output prompt becomes the answer to this query, such as ”the number of active users, traffic load level at specific times, and predicted customer calls next week”. This approach effectively bridges the gap between raw numerical sequences and language-based data representations, facilitating the use of language models for prediction tasks traditionally handled by numerical methods.
@@
Conversely, soft prompts ($P_{\text{soft}}$) introduce trainable embeddings that are optimized during training to influence the model’s prediction subtly [189]. The input for a soft prompt is represented as $[P_{\text{soft}};x_{[1:T]}]$, where $P_{\text{soft}}$ constitutes a series of parameters that are fine-tuned to enhance the predictive capability of the model. This adjustable approach allows the model to internalize and apply nuanced guiding signals without the rigidity of fixed textual cues. Fig. 18 shows an example of using soft prompts in [189]. Upon receiving a time series input, it undergoes tokenization and embedding through a process called patching, complemented by a specialized embedding layer (patch reprogram). Subsequently, all parameters of the pre-trained LLM are frozen, and it requires training solely for the tailored embedding layer to bridge the gap between time series and textual data. When designing soft prompts for time-series prediction using a pre-trained LLM, there are a couple of guiding principles and design choices. Unlike hard prompts, soft prompts require no explicit textual additions to the input data. The general approach involves encoding time series data into a format that the LLM can process, harnessing its underlying capabilities to discern patterns and generate predictions. To utilize soft prompts prediction for telecom efficiently, one might consider the specific characteristics of the telecom time-series data, such as traffic patterns or usage trends, to design the transformation and reprogramming steps that align the time-series data with the model’s language understanding capabilities.

### v2

@@ VI-B LLM-aided Reinforcement Learning for Network Optimization
Reinforcement learning is one of the most important techniques for network optimization. It explores various sequential action combinations, e.g., network resource allocation strategies and signal transmission power level, to maximize the long-term reward, such as higher data rate or lower transmission delay[183]. Many network optimization problems can be transformed into a unified Markov decision process (MDP), and then using reinforcement learning to improve network metrics dynamically. For instance, resource allocation is a very common problem in many telecom scenarios, in which allocation decisions, desired network performance metrics, and network dynamics are usually defined as actions, rewards, and states, respectively[6]. However, it is worth noting that these definitions are usually intuitive and require expert knowledge of reinforcement learning techniques and telecom. Especially, most reward functions are manually designed using trial-and-error approaches, and the algorithm performance is affected by the hyperparameter selection, e.g., learning rate, batch size, and number of hidden layers. Fortunately, LLM techniques provide new opportunities to overcome these bottlenecks. This section will introduce two LLM-aided reinforcement learning techniques: automatic reward function design and verbal reinforcement learning.
@@ VI-B1 Using LLM for reward function design
A recent survey in [27] shows that 92% reinforcement learning researchers use manual trial-and-error reward function design and 89% indicate that the designed rewards lead to unintended behaviour during algorithm training[184]. Such issues become more difficult in complicated telecom scenarios since various network elements are involved, e.g., users with diverse requirements, limited available resources, and dynamic network environments. To this end, LLM shows the capability of developing a universal approach for reward design, which will significantly lower the difficulty of using reinforcement learning. For instance, Song et al. [42] proposed a self-refined LLM for automated reward function design in robotics, achieving a comparable performance as manually designed functions. Kwon et al. [43] applied LLM as a proxy reward function, where the user provides a textual prompt with a few examples or a description of the desired behaviour. In the following, we will introduce how the automatic reward function is designed.
@@
  • A mapping function $\mathcal{M}$ that maps the textual output of LLM to a binary value, e.g., ”good” or ”bad”. This binary value feedback indicates the quality criteria of the generation, and then the LLM easily understands the overall feedback.
Given these definitions, the LLM-aided MDP definition becomes $<S,A,R,T,\mathcal{L},\mathcal{M}>$, where $\mathcal{L}$ is a set of prompts $l_{1}$ to $l_{4}$, and $\mathcal{M}$ is the mapping function. The reward function $\mathcal{F}$ in $R=\mathcal{F}(s,a)$ is defined by $\mathcal{F}:=\mathcal{G}(\mathcal{L},\mathcal{M})$, where $\mathcal{G}$ is the inference of LLMs. $\mathcal{F}:=\mathcal{G}(\mathcal{L},\mathcal{M})$ shows that the design of the reward function $\mathcal{F}$ depends on prompt input $\mathcal{L}$ and the mapping function $\mathcal{M}$. Based on the LLM-aided MDP framework, using LLM for reward function design can be summarized as the following steps:
  • Step 1: Description input. Using language to describe the task, objective, states, and actions. If necessary, providing possible trajectory examples to LLMs. Here an alternative approach is to feed the environment code to the LLM agent, and then using natural language to describe the task, which has been used in [44].
@@
  • Step 4: Evaluation and feedback. Evaluating the reinforcement learning training output and providing feedback to LLMs.
@@ VI-B2 Verbal reinforcement learning via LLM
  • Actor: The actor is built upon an LLM, which is specifically prompted to generate actions, e.g., network control and optimization strategies. Based on short-term and long-term memories, the actor can apply various methods to produce actions, such as CoT [41] and ReAct [186]. These advanced prompt techniques can improve the actor’s capability of reasoning and planning, which can better adapt to the complicated decision-making of network optimization problems.
@@
Compared with conventional reinforcement learning, the LLM-aided verbal learning technique has multiple advantages for telecom optimization: 1) Lowering the difficulty of implementing network optimization. Verbal reinforcement learning avoids the difficulty of tuning hyperparameters like learning rate, batch size, and training frequency. This will significantly lower the difficulty of applying artificial intelligence to network optimization. 2) Allowing for language instructions to guide network optimization policies. Specifically, the LLM-aided system allows for language instructions to guide the agent exploration, which is much more efficient than existing strategies such as $\epsilon$-greedy policy. Experienced network operators can provide language instructions to LLMs directly, and no ML knowledge is required. 3) Reasoning and interpretable explanations for algorithm performance. One crucial advantage of LLM-aided systems is that they provide interpretable explanations of the algorithm and telecom system performance, and these experiences can further help understand network management policies.
Despite the advantages, LLM-aided reinforcement learning is still at a very early stage, and there are very few studies that apply this technique to the telecom field. In addition, specific telecom domain knowledge may be required to let the LLM better understand user demand. Therefore, professional wireless knowledge datasets such as TeleQnA in [125] may be required to fine-tune the LLM.
@@ VI-D LLM-aided Convex Optimization for Telecom
  • Telecom knowledge and formulation templates: Telecom optimization requires professional network knowledge. LLM has learned fundamental knowledge in the pre-training period such as calculating information capacity using Shannon’s formula. However, using state-of-the-art telecom knowledge and formulation templates to fine-tune the LLM can better improve the modelling accuracy. For instance, a dataset named TeleQnA is defined in [125], and it includes nearly 10000 communication field questions from both standards and research articles.
@@
Figure 16: LLM-aided meta-heuristic algorithm generation.
@@ VI-E LLM-based Heuristic Algorithm Design
Heuristic algorithms are very useful techniques for network management and optimization. Specifically, they apply diverse heuristic rules to select near-optimal solutions with low design and computational complexity[192]. Heuristic algorithms are particularly useful for solving optimization problems with integer control variables, which are very frequently formulated in telecom. For instance, the phase-shift optimization of RISs is considered as a very difficult problem with integer control variables and large solution space, and genetic algorithms and particle swarm optimization are used in [193] and [194] to solve this problem. In addition, heuristic algorithms are intuitively compatible with LLMs, since heuristic rules can be easily described by natural language and instructions. For example, swarm-based methods are very widely used heuristic algorithms, e.g., genetic algorithm, particle swarm optimization, and grey wolf optimizer, providing near-optimal solutions by iteratively searching for better solutions. However, the number of these algorithms has grown significantly in the past decade, and selecting the proper algorithm to solve specific network optimization problems has become more difficult. Given the reasoning and understanding capabilities, LLM offers promising changes for selecting and designing novel meta-heuristic algorithms.
@@ VI-F Discussions and Analyses
In addition, several of the above optimization approaches rely on the feedback mechanism, in which the solutions are iteratively improved based on previous answers and environment feedback. For instance, the reward function design is iteratively improved by involving the evaluation results and feedback prompts. Similarly, in verbal reinforcement learning, the LLM agent can adjust the action selections to obtain a higher reward based on environmental feedback. Therefore, the design of these prompts is crucial for improving LLM’s performance, e.g., dedicated feedback and evaluator prompt designs. On the other hand, balancing exploration-exploitation is a common obstacle for many optimization problems. This problem becomes severe when the action space is larger, which is very common in telecom networks. Therefore, how to use the LLM’s self-improvement capability and meanwhile balance the exploration-exploitation is very important.
Finally, note that many optimization problems can be very complicated in wireless networks by involving multiple control variables, network elements and layers. It may require step-by-step problem decomposition, formulating multiple objectives, and alternating optimization, e.g., jointly optimizing the transmission data rate and signal coverage. Handing these optimization problems needs strong planning capabilities, which is still a challenge for current LLM research fields.
@@ VII-A Motivations
Firstly, LLMs provide a universal and generalizable model for telecom network prediction. Given historical data, conventional prediction approaches must train a new model to adapt to incoming target tasks. These methods usually require extensive feature engineering and manual tuning, which can be time-consuming and may not generalize well across different scenarios. By contrast, the versatility of LLMs makes them suitable for processing diverse forms of time-series data, and such adaptability is crucial given the vast volumes of data generated in telecom. Moreover, LLM’s capability to continuously learn and adapt to new data patterns helps mitigate the concept drift problem, ensuring that the models remain relevant and effective over time. As a result, the integration of LLM techniques in time-series prediction offers a promising avenue for developing more robust and generalizable models that can better handle the complexities of data in telecom.
Meanwhile, LLMs have excellent ICL capabilities, which means that they can perform new tasks by leveraging contextual information in demonstrations. In particular, it means that the LLM can directly learn from the provided examples, and map the input-output relationships without extra model training. Such a prediction method is much more efficient than conventional prediction methods. Meanwhile, it is also more accessible since no professional knowledge of model training/fine-tuning is required. In addition, multi-modal LLM-enabled prediction can also be combined with sensing in telecom networks. In particular, multi-modal LLMs can process and integrate information from various data types, such as text, images, audio, and time-series data. In addition, sensing is an important part of envisioned 6G networks, aiming to integrate environmental information into communication networks, e.g., the image captured by street cameras or satellites, 3D LiDAR maps and WiFi sensing. In the context of telecom prediction, multi-modal LLMs can combine sensing data with numerical time-series data to generate more accurate context-aware prediction, which can be particularly useful in 6G.
@@ VII-B Pre-training Foundation Models for Zero-shot Prediction
The pursuit of training a general-purpose foundation model for time-series data is driven by the desire to address the inherent challenges associated with diverse and dynamic data. Traditional time-series methods may struggle to adapt to the non-stationary properties of real-world data, where the statistical characteristics of the series can change over time due to evolving patterns and trends. For instance, the network traffic load level can be affected by many factors, including time, area, environment buildings, service types, etc. It usually requires dedicated model design and training for each target task, and then extracts the underlying patterns from history datasets[197]. By contrast, a general-purpose foundation model aims to overcome these challenges by leveraging the advancements in LLM technologies. The following will first formulate the problem of training a foundational LLM for time-series prediction, and then discuss different tokenization mechanisms and model architectures.
@@ VII-B3 Model architecture
Decoder-only: Even though encoder-decoder models exhibit impressive effectiveness for handling sequences, decoder-only models become more popular in recent years. Like the encoder-decoder architecture, the decoder-only model must first tokenize the raw inputs and then incorporate positional embeddings. The essential difference between encoder-decoder and decoder-only models is that the bidirectional attention is used by encoder [50], which means each token is attending to all other tokens. In contrast, the decoder-only model employs casual attention, where each token cannot attend to tokens after it, but can only look at tokens before it. The causal attention mechanism enhances prediction models because it is well-suited for time-series forecasting tasks. In time-series forecasting, we typically predict future values based on historical data. Causal attention allows each token to consider all preceding tokens, meaning it attends to all events that occurred before the current time frame. Moreover, with these small modifications, the attention score matrices in decoder-only models are triangle matrices, which always have full-column rank, resulting in better expressibility [199]. As shown by Fig.18, TimesFM [200] employs decoder-only architecture to train a time-series prediction model. Unlike traditional LLM techniques, which predict one element at a time, TimesFM is designed to predict extended future sequences in a single step, enhancing accuracy for long-term predictions. This flexibility also extends to inference; given a series, the model can predict its immediate future in fewer steps than a model with equal-length input and output segments would require. Such fast inference could be an appealing feature for telecom applications, because many prediction tasks require rapid response to network dynamics, such as channel state, short-term traffic changes, and indoor user locations. Conventional prediction methods usually take a long training time to adapt to such network environment changes. By contrast, TimesFM has the potential to capture short-term patterns instantly, which aligns with the fast-response requirements of telecom networks.
@@
With a backbone model and an effective tokenization mechanism, one can train a time-series prediction foundation model for telecom with a mixture of different datasets. Meanwhile, understanding the tokenization and model architecture differences is crucial for designing and pre-training a time series LLM for telecom applications. For instance, the tokenization mechanism introduced in the previous Section VII-B2 is very useful for telecom applications, since telecom networks are associated with a large number of network devices and end users, generating a huge number of datasets, such as historical CSI, traffic load level[201], and network performance metrics[202]. Therefore, reducing the number of input tokens can lower the pre-training difficulty of LLMs, especially considering that network edge devices usually have limited computational resources.
@@ VII-C Frozen Pre-trained LLM for Prediction
Rather than developing a specific LLM for prediction, frozen pre-trained LLM refers to approaches that directly adapt a general-domain LLM to prediction tasks. This section delves into using a pre-trained LLM for prediction tasks without the necessity for further fine-tuning. There are two primary approaches: prompting-based and preprocessing-based methods. Specifically, the prompting-based methods include hard and soft prompts. Hard prompts employ rigid and predefined textual structures to present time-series information in a format that is intuitive for the language model. Conversely, soft prompts adopt a more nuanced strategy by integrating trainable embeddings within the input that subtly guide the language model’s predictions. Meanwhile, preprocessing-based methods aim to reformat the time series numerical values into a representation that aligns more seamlessly with LLM’s tokenization process, rather than introducing extra template tokens or trainable embeddings.
@@
Hard prompts ($P_{\text{hard}}$) involve pre-pending a fixed textual instruction and query to the input data sequence, or fit the raw input data into a carefully designed template. In this way, we can transform numerical values into textual contexts that can be processed by pre-trained large language models. By leveraging the impressive generalizability of pre-trained LLMs, hard prompting techniques can yield high prediction accuracy in zero-shot settings. Specifically, the model input for a time-series $x_{[1:T]}$ with a hard prompt is thus formalized as the concatenation $[P_{\text{hard}};x_{[1:T]}]$, which directs the model to generate a prediction in response to the prompt. When designing hard prompts for time-series prediction with language models, the general guideline is to transform numerical data into a format that mimics natural language constructs [203]. This involves two main components: input prompts and output prompts. Input prompts provide historical context and highlight the target time step for prediction, while output prompts focus on the desired prediction value, serving as the ground truth label for training or evaluation. Table X presents several telecom examples of designing hard prompts for specific tasks, such as network traffic load prediction, network user number prediction, and customer service prediction. The process mirrors the source/target structure common in machine translation tasks or can be likened to a question-answering setting, with the context as background information and the question seeking future insights. Then the output prompt becomes the answer to this query, such as ”the number of active users, traffic load level at specific times, and predicted customer calls next week”. According to PromptCast [203], this simple approach achieves comparable or superior prediction accuracy across various datasets, demonstrating its effectiveness on bridging the gap between raw numerical sequences and language-based data representations, further facilitating the use of language models for prediction tasks traditionally handled by numerical methods.
@@
Conversely, soft prompts ($P_{\text{soft}}$) introduce trainable embeddings that are optimized during training to influence the model’s prediction subtly [47]. The input for a soft prompt is represented as $[P_{\text{soft}};x_{[1:T]}]$, where $P_{\text{soft}}$ constitutes a series of parameters that are fine-tuned to enhance the predictive capability of the model. This adjustable approach allows the model to internalize and apply nuanced guiding signals without the rigidity of fixed textual cues. Fig. 19 shows an example of using soft prompts in [47]. TIME-LLM utilizes two types of inputs: a textual description of domain knowledge with some in-context examples, and a time-series input. The textual description is tokenized and processed through the embedding layers of the pre-trained LLM to generate latent representations, termed as prompt embeddings. When a time-series input is received, it is tokenized and embedded through a method called patching, which includes a specialized embedding layer (patch reprogram), resulting in patch embeddings. The pre-trained LLM then takes the concatenated prompt and patch embeddings as inputs and produces outputs via an output projection layer. Throughout this process, all parameters of the pre-trained LLM remain frozen, requiring training only for the custom embedding layer to connect the time series and textual data. When designing soft prompts for time-series prediction using a pre-trained LLM, there are a couple of guiding principles and design choices. Unlike hard prompts, soft prompts require no explicit textual additions to the input data. The general approach involves encoding time series data into a format that the LLM can process, harnessing its underlying capabilities to discern patterns and generate predictions. To utilize soft prompts prediction for telecom efficiently, one might consider the specific characteristics of the telecom time-series data, such as traffic patterns or usage trends, to design the transformation and reprogramming steps that align the time-series data with the model’s language understanding capabilities.